# Biased AI can Influence Political Decision-Making

### v2

@@ Abstract
As modern AI models become integral to everyday tasks, concerns about their inherent biases and their potential impact on human decision-making have emerged. While bias in models are well-documented, less is known about how these biases influence human decisions. This paper presents two interactive experiments investigating the effects of partisan bias in AI language models on political decision-making. Participants interacted freely with either a biased liberal, biased conservative, or unbiased control model while completing political decision-making tasks. We found that participants exposed to politically biased models were significantly more likely to adopt opinions and make decisions aligning with the AI’s bias, regardless of their personal political partisanship. However, we also discovered that prior knowledge about AI could lessen the impact of the bias, highlighting the possible importance of AI education for robust bias mitigation. Our findings not only highlight the critical effects of interacting with biased AI and its ability to impact public discourse and political conduct, but also highlights potential techniques for mitigating these risks in the future.
@@
In recent years, the rapid advancements in modern AI language models have catapulted them to the forefront of our daily interactions, resulting in a fundamental change in how we communicate, gather information, and form opinions. From political news summarization [Hu2023BadAG] to the use of language models for fake news detection [news_summary], AI models are increasingly becoming seamless tools in our daily lives. However, as these models proliferate, concerns have emerged regarding their inherent biases and propensity to generate false information, raising critical ethical and legal questions about their impact on human cognition and decision-making [Elsafoury2023SystematicOS, Li2023TheDS, forbes_bias_llm, nyt_bias, pnas.2313790120].
Modern AI language models have repeatedly been shown to exhibit inherent specific behavioral biases such as social bias [wan_yuxan, Fang2023BiasOA], partisan bias [röttger2024politicalcompassspinningarrow, feng2023pretrainingdatalanguagemodels], and other demographic representation bias [kirk2021biasoutoftheboxempiricalanalysis, racist_bias_dialect]. This bias has been shown to permeate many different stages of these models, including training data [zhao-etal-2019-gender, 10.1145/3442188.3445922], word embeddings [zhao-etal-2019-gender, 10.5555/3157382.3157584, nissim-etal-2020-fair], model architecture [blodgett-etal-2020-language, hovy_bias], and output [BaumSethD2024MASv, Mittermaier2023BiasIA]. Moreover, it has been shown that bias can be easily introduced in a model through methods as simple as the phrasing of the language model input prompts or instructions [wan_yuxan, lin-ng-2023-mind, Cantini2024AreLL]. Addressing bias in models is a complex challenge, and developing efficient methods to mitigate it continues to be a focus of ongoing research [Mittermaier2023BiasIA, oconnor_gender_bias, Srivastava2024DeamplifyingBF]. Despite the well-documented presence of bias in language models, the critical question of whether these biases have a measurable influence on human decision-making—and under what circumstances this influence is heightened or diminished—remains less clear.
@@
Research on the effects of biased AI language models on attitudes and behavior is limited, has yielded unclear results, and has mainly focused on inconsequential decisions. For instance, some recent studies find that biased AI-generated information can influence decisions in areas such as medical classifications and educational hiring [Wambsganss2023UnravelingDG, bias_politics_gpt2, vicente_bias_medical]; however, these findings are based on static AI-generated content and often involve fictional or impersonal tasks, which may increase participants’ susceptibility to influence by not engaging their personal values. Similarly, studies examining AI-generated autocomplete suggestions involve more dynamic interactions between language models and users, but their results are mixed, with some showing an influence and others not [Wambsganss2023UnravelingDG, Jakesch2023CoWritingWO].
@@
In contrast, a robust body of research has shown that interactions with biases in traditional forms of communication does influence human decision-making [bias_media_meta]. For example, research indicates that humans are affected when engaging with biased individuals [bias_media_meta], biased print media [news_affect_cancer], and consuming biased political news outlets [AGGARWAL2020100025, druckman_media_bias, broockman_bias_media].
@@
To bridge this gap, we conducted a series of experiments to evaluate the impact of biased AI language models on human decision-making in a more typical setting, using dynamic chatbox interactions, with tasks centered on personal opinions and decisions. Specifically, we examine the impact of model bias on political decision making, which has not been previously studied, by deploying two sets of experiments in which individuals who identified themselves as Democrats or Republicans were asked to make decisions about U.S. political topics after discussing these topics with an AI language model. For this paper, we focus on language model behavioral bias, which we define as the variations in generated text, where the model’s responses—such as recognizing, rejecting, or reinforcing stereotypes—change based solely on the social group mentioned in the prompt [Kumar2024DecodingBA]. The type of model bias we examine is partisan bias, which we define as the the tendency of political partisans to process information and make judgments in a way that favors their own party [iyengar2019origins, NBERw19080].
@@
In the first experiment, participants were asked to form unidimensional, pro- or anti-, opinions on a number of political topics that they were unlikely to have prior opinions on. In the second experiment, decisions were open-ended, as participants were asked to distribute funds to four different sectors of government (K-12 Education, Welfare, Safety, and Veterans). In both cases, participants were randomly and unknowingly assigned to interact with biased liberal, biased conservative, and unbiased control language models to evaluate the effects of these interactions. We focus on partisan bias due to its known prevalence in current state-of-the-art models [röttger2024politicalcompassspinningarrow, feng2023pretrainingdatalanguagemodels], the high-level of public concern about such bias, and feasibility due to its polarized, salient nature.
These experiments found clear evidence that participants were swayed by partisan bias in AI language models, regardless of their prior political beliefs or whether those beliefs aligned with the bias of the language models. Even more so, we found that even correct detection of bias in the model did not significantly reduce the effect. However, prior self-reported knowledge of AI technology did slightly reduce the effects of the bias models. By focusing on partisan bias, we aim to shed light on the potential consequences and ethical considerations associated with the use of biased AI language models in shaping public discourse, opinion formation, and political behavior. Furthermore, this study is among the first to determine whether dynamic interactions with biased AI language models directly influence human decision-making and opinions, as well as focusing on decisions and opinions that are often driven by personal values and identity.
1 Results
Each participant was given two tasks to complete: the Topic Opinion Task and the Budget Allocation Task. In the Topic Opinion Task, participants were asked to indicate their baseline knowledge and opinions about two relatively obscure political topics — one typically supported by liberals and the other by conservatives. They were then instructed to interact with an AI language model (such as ChatGPT) and were randomly assigned to either a liberal-biased language model, a conservative-biased language model, or a control model. Following this interaction, they were again asked to indicate their level of knowledge and opinion. For the topics used see Table 3.
For the Budget Allocation Task, participants were asked to imagine themselves as the mayor of a city with leftover funds to distribute among four government funding areas, two of which were generally prioritized by liberals and two by conservatives. Participants made their initial allocation decisions and submitted them to the randomly assigned AI language model for feedback. After receiving feedback, participants were encouraged to interact with the AI through a chat interface to ask follow-up questions and seek further clarification. Following these interactions, participants submitted their final allocation. In both tasks, participants were required to have at least three and up to twenty interactions with the model.
We created each treatment condition by instructing the model to respond with the designated bias. These instructions were written in the back-end code and were unknown to the participant. The control was instructed to respond as a “neutral” American. The evaluation of partisan biases in the treatment models aligned with expectations based on the Political Compass Test, a 62-question assessment that measures political leanings across two dimensions: economic and social (see Figure 3). Participants were not informed of the underlying purpose of the study prior to completing the tasks; they were simply told that they would be interacting with an AI language model to complete each tasks. To determine how the AI language model affected the degree of change in opinion due to model bias, we compared the pre- and post-interaction opinion scores of the biased AI language models to the control AI language models, thus controlling for the effect of interacting with an AI language model.
@@ 1 Results
Interaction with Biased AI Affects Political Opinions
@@
In the Topic Opinion Task, we found that participants who interacted with biased language models were more likely to change opinions in the direction of the bias of the language model compared to those who interacted with the neutral AI language model, even if it was opposite to what their beliefs were likely to be, based on their stated political affiliation. We found that on topics typically aligned with conservative views, Democrats who were exposed to liberal-biased models significantly reduced support for conservative topics after interactions compared to those exposed to the neutral models (coefficient-value = -0.85, t = -2.38, p-value = 0.02), and those exposed to conservative-biased models significantly increased support for conservative topics compared to those exposed to the neutral models (coefficient-value = 0.98, t = 2.71, p-value = .007). Similarly, Republican participants who interacted with the liberal-biased model had reduced support for the conservative topic compared to the Republicans who interacted with the neutral model (coefficient-value = -0.79, t = -2.16, p-value = .03). However, Republican participants exposed to the conservative-bias model did not have a statistically significant difference in opinions compared to those exposed to the neutral model. This is likely representing a ceiling effect, as these participants already agreed strongly with the model’s bias and therefore had little room to further increase their support. See Table 1 (top) for full results.
@@
For topics aligned with liberal preferences, we found that both Republicans and Democrats who were exposed to conservative AI models had a statistically significant decrease in support for the topic compared to those who were exposed to the neutral model (coefficient value = 1.42, t = 3.91, p-value < 0.001 and coefficient-value = 1.44, t = 3.82, p-value < 0.001, respectively). However, exposure to a liberal model did not have an effect of increasing support for the topics with either group compared to the neutral model. See Table 1 (bottom) for full results. We also conducted the same analysis subsetting only to participants who indicated no prior knowledge of the topics and the results remain unchanged, indicating that interacting with biased AI language models affects opinion formation as well (see Section E.2 for details).
We note that for the liberal aligned topics, the neutral AI language model led to an unexpected shift in the post-interaction baseline for both Democrats and Republicans towards a liberal position. Thus, there was a ceiling effect, whereby exposure to a liberal biased AI language model could not further shift an already moderate liberal opinion when interacting with the neutral model. One possible explanation for the liberal shift from the control model is that partisan respondents do not exhibit expected consistency in ideological beliefs on low salience issues with multiple dimensions [lenz2012follow, doi:10.1086/700005]. All issues have multiple dimensions and partisan alignment may depend on which dimension is more salient. Elite signaling informs partisans of the issues they should oppose or support, but that is absent for the low salience issues we chose. For more discussion, see Section E.1.
@@
Qualitatively, in this task, participants often treated the model like a traditional search engine, with of initial interactions involving queries such as “What is <topic>?”. Common follow-up questions included “What are the pros/cons of <topic>?” or more specific inquiries like “How many states offer covenant marriages?” or “Does the US practice unilateralism in foreign relations?”. Although participants mainly sought information from the model in the form of questions, we did find that asked the model for its opinion on the topic. Another used some form of conversational language such as “hello”, “good afternoon”, “I see”, or “thank you”, suggesting that they found it to be more human-like than a simple search engine. Some even seemed to argue with the model, when it was not aligned with their views, or find comradery when it did. For examples of these conversations, see Section E.5.
| Participant Partisanship | Branch | ANOVA (p-value) | Dunnett Test | Dunnett (p-value) |
| Democrat | Safety | <0.01 | Liberal vs. Control | <0.01 |
| Conserv. vs. Control | 0.13 | |||
| Veterans | <0.01 | Liberal vs. Control | 0.01 | |
| Conserv. vs. Control | <0.01 | |||
| Education | <0.01 | Liberal vs. Control | 0.03 | |
| Conserv. vs. Control | <0.01 | |||
| Welfare | <0.01 | Liberal vs. Control | 0.01 | |
| Conserv. vs. Control | 0.08 | |||
| Republican | Safety | <0.01 | Liberal vs. Control | <0.01 |
| Conserv. vs. Control | <0.01 | |||
| Veterans | <0.01 | Liberal vs. Control | 0.60 | |
| Conserv. vs. Control | 0.03 | |||
| Education | <0.01 | Liberal vs. Control | 0.03 | |
| Conserv. vs. Control | <0.01 | |||
| Welfare | <0.01 | Liberal vs. Control | 0.06 | |
| Conserv. vs. Control | 0.03 |
@@
Interaction with Biased AI Affects Political Decision-Making
@@
In the Budget Allocation Task, we found strong evidence that participants who interacted with biased language models were more likely to change their proposed budget allocation to be aligned with the bias of the AI language model compared to those who interacted with the neutral model, again even when the bias was opposed to their stated political values. We found that the change in budget allocation towards the biases of the models compared to the control model for all participants, regardless of personal ideology, was highly statistically significant with , see Table 2.
@@
Figure 1 shows the average change in allocation in each of the experimental conditions and control for both groups of participants. We found that the largest average change ( confidence interval) was demonstrated for Democrat participants when exposed to the conservative AI model with average changes of () for Education, () for Welfare, () for Safety and () for Veterans. Similarly, the largest change in allocation for Republican participants was when they are exposed to the liberal AI model with average changes ( confidence interval) of () for Education, () for Welfare, () for Safety, and () for Veterans. This task showed that interacting and collaborating with biased AI had strong effects on the change in outcome and final allocation of the budgets proposed.
@@
Qualitatively, we found participants in this task were more likely to interact with the model conversationally and collaboratively, with of the participants asking for the model’s opinion on the allocation. However, only asked the model information-based questions such as “Do any of these four funding areas receive federal or state funding” or “Is there a correlation between public safety investment and lower crime rates?”. The interactions were more focused on the collaboration with the model and the expression of opinions, as seen in the example conversation in the Section E.5.
@@
We hypothesized that prior general knowledge of AI might reduce the effects of interaction with the biased AI language model, as individuals with some understanding of AI might be more aware of the potential biases and reliability limitations of large language models. To test this hypothesis, we included a binary indicator of self-reported prior AI knowledge compared to the general population (“more knowledge” or “less knowledge”) as a control variable in our ordinal regression and ANOVA test for the Topic Opinion Task and Budget Allocation Task respectively. We note that only 32% (n=) of Democrats and 47% (n=71) of Republicans indicated having more AI knowledge. Even with this low power, we did find some evidence to support this hypothesis. Specifically, for the conservatively supported topics in the Topic Opinion Task with Democrat participants, we found that having prior knowledge of AI significantly reduced the influence of the biased AI interactions compared to those who had less prior knowledge of AI (coefficient value = -0.79, t value = -2.51, p value =.01). For the Budget Allocation Task, we found significant differences at the level in the allocation between participants who reported more or less AI knowledge in the Veterans funding allocation for Democrat participants (p-value = 0.09) and Safety funding allocation for Republican participants (p-value = 0.08). See Section E.3 for the full results. These findings give some indication that prior knowledge of AI might reduce the effects of biased AI language models.
A second hypothesis, which has been supported in the context of traditional media, suggests that participants who recognize a news source’s bias are less influenced by the bias [kroon_etal]. Thus, we aimed to test whether awareness of bias in the model responses would also reduce the effect of the bias on participants. To do this, we added a binary indicator bias detection, where “correct” detection meant that a participant was in a bias condition and correctly identified the model as biased by answering “likely yes” or “definitely yes” when asked if the model was biased. The answer was labeled as “incorrect” if it was “likely no” or “definitely no”. Since we are interested in Type 2 errors only, we used all participants in the control condition, regardless of their bias detection. We note that (n=) of Democrat and (n=) of Republicans in a bias conditions correctly identified bias in the model. Again, we included this binary bias detection variable as a control variable in our ordinal regression and ANOVA test for the Topic Opinion Task and the Budget Allocation Task respectively. For the Topic Opinion Task and Budget Allocation Task we did not find the coefficient significant in any condition. See Section E.3 for the full results. This suggests that participants who were aware that the AI model was biased were affected similarly to those who were not aware of the bias.
@@
The collaborative nature of the Budget Allocation Task provided a unique opportunity to explore the persuasion techniques used across experimental conditions, offering valuable insights for model bias mitigation strategies. To analyze the conversations, we annotated them using the latest GPT-4 model [gpt4], employing a list of persuasion techniques compiled from a meta-analysis of persuasive strategies [Piskorski2023SemEval2023T3]. To ensure quality, we conducted a human evaluation of of the model’s annotations, achieving accuracy. Our analysis found no significant differences in the distribution of persuasion techniques between the experimental conditions and the control group, as determined by a Chi-square test with Monte Carlo correction ( = 24.5, p = .07). Across all three conditions, the most frequently used techniques used by the AI language models were “Appeal to Values,” “Consequential Oversimplification,” “Appeal to Authority,” and “Repetition” (see Figure 2(a)).

### v3

# Biased LLMs can Influence Political Decision-Making

@@ Abstract
As modern large language models (LLMs) become integral to everyday tasks, concerns about their inherent biases and their potential impact on human decision-making have emerged. While bias in models are well-documented, less is known about how these biases influence human decisions. This paper presents two interactive experiments investigating the effects of partisan bias in LLMs on political opinions and decision-making. Participants interacted freely with either a biased liberal, biased conservative, or unbiased control model while completing these tasks. We found that participants exposed to partisan biased models were significantly more likely to adopt opinions and make decisions which matched the LLM’s bias. Even more surprising, this influence was seen when the model bias and personal political partisanship of the participant were opposite. However, we also discovered that prior knowledge of AI was weakly correlated with a reduction of the impact of the bias, highlighting the possible importance of AI education for robust mitigation of bias effects. Our findings not only highlight the critical effects of interacting with biased LLMs and its ability to impact public discourse and political conduct, but also highlights potential techniques for mitigating these risks in the future.
Biased LLMs can Influence Political Decision-Making
Jillian Fisher 1, Shangbin Feng 2, Robert Aron 3, Thomas Richardson 1, Yejin Choi 4, Daniel W. Fisher 5, Jennifer Pan 6, Yulia Tsvetkov 2, Katharina Reinecke 2 1 Department of Statistics, University of Washington, 2Department of Computer Science, University of Washington, 3Dallas, Texas, 4Department of Computer Science, Stanford University 5Psychiatry and Behavioral Science, University of Washington, 6Department of Communication, Stanford University Correspondence: jrfish@uw.edu
1 Introduction
@@
In recent years, the rapid advancements in modern large language models (LLMs) have catapulted them to the forefront of our daily interactions, resulting in a fundamental change in how we communicate, gather information, and form opinions. From political news summarization Hu et al. (2023) to the use of language models for fake news detection Zhang et al. (2024), LLMs are becoming seamlessly integrated into our daily lives. However, as these models proliferate, concerns have emerged regarding their inherent biases and propensity to generate false information, raising critical ethical and legal questions about their impact on human cognition and decision-making Elsafoury et al. (2022); Li (2023); Knapton (2023); Metz (2023); Acerbi and Stubbersfield (2023).
@@
Research on the effects of biased LLMs on attitudes and behavior is limited or has yielded unclear results. For instance, some recent studies find that biased LLM-generated information can influence decisions in areas such as medical classifications and educational hiring Wambsganss et al. (2023); Liu et al. (2022); Vicente and Helena (2023); however, these findings are based on static LLM-generated content and often involve fictional or impersonal tasks, which may increase participants’ susceptibility to influence by not engaging their personal values. Similarly, studies examining LLM-generated autocomplete suggestions involve more dynamic interactions between language models and users, but their results are mixed, with some showing an influence and others not Wambsganss et al. (2023); Jakesch et al. (2023).
@@
In contrast, a robust body of research has shown that long-term interactions with biases in traditional forms of communication does influence human decision-making DellaVigna and Kaplan (2008). For example, research indicates that humans are affected when engaging with biased individuals DellaVigna and Kaplan (2008), biased print media Jensen et al. (2014), and consuming biased political news outlets Aggarwal et al. (2020); Druckman and Parkin (2005); Broockman and Kalla (2024). However, LLMs introduce new complex dynamics, particularly due to their being perceived as both authoritative and objective while simultaneously facing widespread global distrust from users Gillespie et al. (2023); University (2024). These unique factors may amplify or diminish the effect of bias in ways different from traditional sources such as media, warranting a specific investigation.
@@
To bridge this gap, we conducted a series of experiments to evaluate the impact of biased LLMs on human decision-making in a more typical setting, using dynamic chatbox interactions, with tasks centered on personal opinions and decisions. Specifically, we examine the impact of model bias on political decision making, which has not been previously studied, by deploying two sets of experiments in which individuals who identified themselves as Democrats or Republicans were asked to make decisions about U.S. political topics after discussing these topics with an LLM. For this paper, we focus on language model behavioral bias, which we define as the variations in generated text, where the model’s responses—such as recognizing, rejecting, or reinforcing stereotypes—change based solely on the social group mentioned in the prompt Kumar et al. (2024). The type of model bias we examine is partisan bias, which we define as the tendency of political partisans to process information and make judgments in a way that favors their own party Iyengar et al. (2019); Bullock et al. (2015).
@@
In the first experiment, participants formed unidimensional pro- or anti- opinions on unfamiliar political topics. In the second, they were asked to allocate funds across four government sectors. In both, participants unknowingly interacted with either a liberally biased, conservatively biased, or neutral LLM to assess the effects of partisan bias. We focus on partisan bias due to its prevalence in state-of-the-art models Röttger et al. (2024); Feng et al. (2023), public concern, and its polarized, salient nature. See Figure 1 for an overview of our experimental design.
Results showed that LLM bias influenced participants’ opinions and decisions, regardless of their prior beliefs or alignment with the model’s bias. Surprisingly, even those with opposing political views shifted toward the model’s stance, challenging research suggesting resistance to belief change in short-term interactions Nyhan and Reifler (2010); Lord et al. (1979); Ahluwalia (2000). Notably, recognizing bias in the generations did not reduce its impact, though self-reported AI knowledge slightly mitigated it. By examining partisan bias, this study highlights ethical concerns surrounding biased LLMs in public discourse and is among the first to explore how dynamic interactions with biased models shape human decisions and values.
2 Methods
Each participant completed two tasks: the Topic Opinion Task and the Budget Allocation Task. Both followed a similar structure—a pre-survey, followed by interaction with an LLM via chatbox, and a post-survey. During the interaction, participants engaged freely with an LLM but were unknowingly assigned to either a liberal-biased, conservative-biased, or control model. Full details of our study design can be found in Appendix B.
@@ 2 Methods
Participants
We recruited participants via Prolific Prolific , requiring them to be U.S. citizens over 18, proficient in English, and self-identified as either Republican or Democrat. There were no exclusion criteria. A pilot study ( n=) informed our sample size calculation via simulation power analysis (, ), resulting in n= per political group (total N=) to detect a medium-to-small effect. One participant was removed for inappropriate LLM interaction, leaving N= ( female, male; mean age , SD ). Republicans (n=) and Democrats (n=) were balanced by design. Participants were compensated at /hour. Full demographics are in Section A.3. The study was deemed exempt by a University of Washington IRB; ethical considerations are detailed in Appendix D.
Experimental Setup
Before experimentation, participants were asked to sign an informed consent. Participants were only told they would be interacting with AI language models to complete tasks, but no mention of biased AI was included. Participants were first asked demographic questions including their age, gender, race and ethnicity, their highest level of education, income, and partisan affiliation. Then, participants were asked to complete two tasks, following a consistent three-stage design: an initial choice section where their views on the topic were measured; interaction with an AI language model, where they gathered more information on the topic via typed conversation with the AI language model in a chatbox; and a post-choice section where they were again asked the same questions as the pre-choice section to measure how their opinions had changed. See Section A.1 for experimental overview.
We employed a experimental design, featuring three experimental factors (AI liberal bias, AI conservative bias, AI neutral) and two participant factors (Republican and Democrat participants). After consent and initial data gathering, participants were randomly assigned to an experimental condition (liberal biased AI, conservative biased AI, or neutral AI), an order of the tasks (Topic Opinion Task, and Budget Allocation Task), order of topics in the Topic Opinion Task (liberal support topic and conservative support topic), and specific topic for the Topic Opinion Task (assigned one of the two options per topic type in Table 13). Participants were not informed in any way as to whether the AI language model was biased or neutral. After completion of both tasks, we asked a series of follow-up questions related to the participants’ experience with the AI language model and their overall level of AI knowledge, in general. Finally, we debriefed the participant on the true nature of the study, including the potential bias of the AI, and gave them an option to opt out of the study. No participant chose to opt out of the study.
Experimental Setup: Topic Opinion Task
In the Topic Opinion Task, participants first reported their baseline knowledge and opinions on two relatively obscure political topics—one typically supported by liberals and the other by conservatives. They then freely interacted with an LLM to learn more about the topic before reassessing their knowledge and opinions. Again, the participant was unaware of the potential partisan leaning of the model they were interacting with. Using lesser-known topics helped minimize prior biases Taber and Lodge (2006) and better modeled real-world LLM interactions where users seek information on unfamiliar issues. The selected topics were multifamily housing and the Lacey Act of 1900 (liberal-supported) and international unilateralism and covenant marriages (conservative-supported). Further details on topic selection are in Section B.1.
Experimental Setup: Budget Allocation Task
Inspired by negotiation tasks in group decision theory, particularly the Legislative Task Mennecke et al. (2000); He et al. (2017), the Budget Allocation Task required participants to act as a city mayor distributing remaining government funds among four entities: Public Safety, Education, Veteran Services, and Welfare. These categories were chosen to reflect issues that elicit differing funding priorities among conservatives and liberals (see Section B.2). Participants made their initial allocation decisions and submitted them to the randomly assigned LLM for feedback. After receiving feedback, participants were encouraged to interact with the LLM through a chat interface to ask follow-up questions and seek further clarification. Following these interactions, participants submitted their final allocation. In both tasks, participants were required to have at least three and up to twenty interactions with the model, where an interaction was any question or response sent to the LLM.
Treatment Condition
Participants were randomly assigned to one of three experimental conditions, which remained consistent across all tasks: interaction with a liberal-biased, conservative-biased, or neutral (control) language model. To create these models, we used GPT-3.5-turbo OpenAI (2023) and introduced prefixed identifiers to each input. Using a prefix was preferred over fine-tuning, as it avoided the need for extensive training data while leveraging the fluency of pre-trained models. Participants were not informed of the addition of prefix identifiers prior to treatment.
To validate bias, we followed Feng et al. (2023) by administering the Political Compass Test (PCT) pct , which positions models on economic and social axes. Our results (see Section A.4) confirmed that the liberal-biased model exhibited liberal views, the conservative model leaned conservative, and the neutral model remained centrist, refusing to take a stance on of PCT questions, compared to and for the conservative and liberal models, respectively. Beyond basic prefixing, we decided to further reinforce bias behavior of the LLM by explicitly instructing models on specific topics. For instance, for a liberal-biased response on "covenant marriage," we used the following prefix:
“Respond as a radical left U.S. Democrat. As such, you are not supportive of covenant marriages. <participant input>”
This approach ensured consistent bias across interactions (see Section A.4 for exact prompts).
Analysis
For the Topic Opinion Task, to measure the effect of LLM interactions on opinion shifts, we analyzed the change in opinion before and after participants engaged with the model. We used ordinal logistic regression separately for Republicans and Democrats, modeling opinion change as a function of exposure to liberal or conservative bias, using the following equations,
where represents the difference between post- and pre-interaction responses on a 7-point Likert scale. The magnitude reflects the extent of change and sign indicates direction (negative for liberal shifts, positive for conservative shifts). We tested the significance of bias effects () using t-tests () and extended the model to assess prior knowledge and bias detection . However, since these secondary analyses were not randomized, they provide correlational rather than causal insights.
For the Budget Allocation Task, we examined shifts in budget allocations for the four government areas, using ANOVA to assess changes in allocation (post-pre) per area. We used the same equation above eq. 1, with only a change in . Significant effects were followed by Dunnett post-hoc tests comparing control and bias experimental groups (). As with opinion shifts, we explored the effects of prior knowledge and bias detection , though these findings remain exploratory due to the lack of randomization.
For both the Budget Allocation Task and Topic Opinion Task, we ran a seperate analysis including each demographic variable (see Section B.3 for a list), however, we found no significant changes to the model. Therefore, we did not include any moderating variables related to the differences between the individual participants.
3 Results
Interaction with Biased LLMs Affects Political Opinions
@@ 3 Results
In the Topic Opinion Task, we found that participants who interacted with biased language models were more likely to change opinions in the direction of the bias of the model compared to those who interacted with the neutral model, even if it was opposite to what their beliefs were likely to be, based on their stated political affiliation. We found that on topics typically aligned with conservative views, Democrats who were exposed to liberal-biased models significantly reduced support for conservative topics after interactions compared to those exposed to the neutral models (coefficient-value = -0.85, t = -2.38, p-value = 0.02), and those exposed to conservative-biased models significantly increased support for conservative topics compared to those exposed to the neutral models (coefficient-value = 0.98, t = 2.71, p-value = .007). Similarly, Republican participants who interacted with the liberal-biased model had reduced support for the conservative topic compared to the Republicans who interacted with the neutral model (coefficient-value = -0.79, t = -2.16, p-value = .03). However, Republican participants exposed to the conservative-bias model did not have a statistically significant difference in opinions compared to those exposed to the neutral model. This is likely representing a ceiling effect, as these participants already agreed strongly with the model’s bias and therefore had little room to further increase their support. See Table 1 (top) for full results.
| Participant Partisanship | Branch | Dunnett Test | Dunnett (p-value) |
| Democrat | Safety | Liberal | <0.01 |
| Conserv. | 0.13 | ||
| Veterans | Liberal | 0.01 | |
| Conserv. | <0.01 | ||
| Education | Liberal | 0.03 | |
| Conserv. | <0.01 | ||
| Welfare | Liberal | 0.01 | |
| Conserv. | 0.08 | ||
| Republican | Safety | Liberal | <0.01 |
| Conserv. | <0.01 | ||
| Veterans | Liberal | 0.60 | |
| Conserv. | 0.03 | ||
| Education | Liberal | 0.03 | |
| Conserv. | <0.01 | ||
| Welfare | Liberal | 0.06 | |
| Conserv. | 0.03 |
@@
For topics aligned with liberal preferences, we found that both Republicans and Democrats who were exposed to the conservative model had a statistically significant decrease in support for the topic compared to those who were exposed to the neutral model (coefficient-value = 1.44, t = 3.82, p-value < 0.001 and coefficient value = 1.42, t = 3.91, p-value < 0.001, respectively). However, exposure to a liberal model did not have an effect of increasing support for the topics with either group compared to the neutral model. See Table 1 (bottom) for full results.
We also conducted the same analysis subsetting only to participants who indicated no prior knowledge of the topics and the results remain unchanged, indicating that interacting with biased LLMs affects opinion formation as well (see Section E.2 for details).
Interestingly, we did notice that for liberal-aligned topics, the neutral LLM unexpectedly shifted both Democrats and Republicans toward a more liberal stance, creating a ceiling effect where the liberal-biased model had no further impact. This may stem from partisan inconsistency on low-salience, multi-dimensional issues, where alignment depends on which aspect is most salient. Without elite signaling to guide positions, partisans may deviate from expected ideological patterns Lenz (2012); Freeder et al. (2019). See Section E.1 for further discussion.
@@
Qualitatively, participants largely interacted with the model like a search engine during this task, with of initial queries asking, “What is <topic>?” Common follow-ups included “What are the pros/cons of <topic>?” or specific factual questions like “How many states offer covenant marriages?”. Only about sought the model’s opinion, while used conversational language (e.g., “hello,” “thank you”), suggesting they perceived it as somewhat human-like. Some even argued with the model when it contradicted their views or found camaraderie when it aligned. This qualitative analysis was conducted manually; see Section E.5 for details.
@@
Interaction with Biased LLMs Affects Political Decision-Making
@@
In the Budget Allocation Task, we found strong evidence that participants who interacted with biased language models were more likely to change their proposed budget allocation to be aligned with the bias of the model compared to those who interacted with the neutral model, again even when the bias was opposed to their stated political values. We found that the change in budget allocation towards the biases of the models compared to the control model for all participants, regardless of personal ideology, was highly statistically significant with , see Table 2.
@@
Figure 2 shows the average change in allocation in each of the experimental conditions and control for both groups of participants. We found that the largest average change ( confidence interval) was demonstrated for Democrat participants when exposed to the conservative LLMs with average changes of () for Education, () for Welfare, () for Safety and () for Veterans. Similarly, the largest change in allocation for Republican participants was when they are exposed to the liberal LLMs with average changes ( confidence interval) of () for Education, () for Welfare, () for Safety, and () for Veterans. This task showed that interacting and collaborating with biased LLMs had strong effects on the change in outcome and final allocation of the budgets proposed.
@@
Compared to the Topic Opinion Task, participants in this task engaged with the model more conversationally and collaboratively, with asking for its opinion on budget allocation. In contrast, only sought factual information, posing questions like “Do these funding areas receive federal or state funding?” or “Is there a correlation between public safety investment and lower crime rates?” Overall, interactions emphasized collaboration and opinion exchange rather than information retrieval (see Section E.5 for examples).
@@
We hypothesized that prior AI knowledge might mitigate the influence of biased LLM interactions, as individuals aware of AI’s limitations may be more cautious of its biases. To test this, we included a binary indicator of self-reported AI knowledge (“more” vs. “less” than the general population) as a control variable in our ordinal regression and ANOVA for the Topic Opinion Task and Budget Allocation Task, respectively. However, since this variable was not randomized, our findings are correlational rather than causal. Also, only of Democrats () and of Republicans (n=) reported having more AI knowledge, limiting statistical power. Despite this, we found some evidence supporting our hypothesis. Among Democrats in the Topic Opinion Task, prior AI knowledge significantly reduced the effect of biased interactions on conservatively supported topics (coefficient value = -0.79, t = -2.51, p value =.01). In the Budget Allocation Task, we observed marginally significant differences () in Veterans funding allocation for Democrats (p = .09) and Safety funding allocation for Republicans (p = .08) based on AI knowledge. These results suggest that prior AI knowledge may help mitigate bias effects. However, given the lack of randomization and small sample size, these findings are hypothesis-generating rather than conclusive, warranting further investigation.
A second hypothesis, supported in traditional media studies, suggests that recognizing bias reduces its influence Kroon et al. (2022). We tested whether this applies to LLM-generated content by introducing a binary bias detection variable. Participants in a biased condition were classified as having “correctly” detected bias if they answered “likely yes” or “definitely yes” when asked if the model was biased; responses of “likely no” or “definitely no” were classified as “incorrect.” Since we are interested in Type 2 errors only, we used all participants in the control condition, regardless of their bias detection. Overall, (n=) of Democrats and (n=) of Republicans in a bias conditions correctly identified bias in the model. Again, we included this binary variable as a control in our ordinal regression and ANOVA for the Topic Opinion Task and Budget Allocation Task, respectively. However, as bias detection is a post-treatment variable, it cannot be used as a mediator without potential bias Montgomery et al. (2018). Nonetheless, we include this analysis to align with prior media bias research Chiang and Knight (2011); Han et al. (2022). We found no significant effect of bias detection in any condition for either task (see Section E.3 for full results). This suggests that participants who recognized the LLMs bias were influenced similarly to those who did not.
@@
The collaborative nature of the Budget Allocation Task provided a unique opportunity to explore the persuasion techniques used across experimental conditions, offering valuable insights for model bias mitigation strategies. To analyze the conversations, we annotated them using the latest GPT-4 model OpenAI (2024), employing a list of persuasion techniques compiled from a meta-analysis of persuasive strategies Piskorski et al. (2023b). To ensure quality, we conducted a human evaluation of of the model’s annotations, achieving accuracy. Our analysis found no significant differences in the distribution of persuasion techniques between the experimental conditions and the control group, as determined by a Chi-square test with Monte Carlo correction ( = 24.5, p = .07). Across all three conditions, the most frequently used techniques used by the LLMs were “Appeal to Values,” “Consequential Oversimplification,” “Appeal to Authority,” and “Repetition” (see Figure 3 - left).